OpenStudy (anonymous):

Let two positive intigers be such that their sum is 9 and the sum of their fourth powers is 2417. Find the two numbers?

OpenStudy (anonymous):

@Arnab09 @yash2651995

Parth (parthkohli):

\( \color{Black}{\Rightarrow a + b = 9 }\) \( \color{Black}{\Rightarrow a^4 + b^4 = 2417 }\) Although you may attempt a trial and error here.

OpenStudy (lgbasallote):

this doesnt seem easy lol

Parth (parthkohli):

8,1 ; 7,2 ; 3,6 ; 4,5

Parth (parthkohli):

Check if any of em satisfy the second equation.

OpenStudy (lgbasallote):

saw one :D

Parth (parthkohli):

Yes... I got it ;D

Parth (parthkohli):

Go put these in your calculator. \( \color{Black}{\Rightarrow 8^4 + 1^4 }\) \( \color{Black}{\Rightarrow 7^4 + 2^4 }\) \( \color{Black}{\Rightarrow 6^4 + 3^4 }\) \( \color{Black}{\Rightarrow 5^4 + 4^4 }\)

OpenStudy (yash2651995):

\[(a+b)^{4}= a ^{4}+b ^{4}+4(ab)(a+b)+6(ab)^{2}\] find (ab) from here .. then put it in \[(a-b)^{2}=(a+b)^{2}-4(ab)\] then find a-b use a+b=9 and a-b= ..what ever you get.. to find a and b

OpenStudy (yash2651995):

ab will be positive..^

OpenStudy (shubhamsrg):

your formulla is wrong @yash2651995 please recheck

OpenStudy (yash2651995):

ya.. let me correct it..

OpenStudy (shubhamsrg):

formulla is (a+b)^4 = a^4 + 4a^3 b + 6a^2 b^2 + 4ab^3 + b^4 => a^4 + b^4 + ab ( 4a^2 + 4b^2 + 6ab)

OpenStudy (shubhamsrg):

and a^2+ b^2 = 81 -2ab now you can make this subsn and this should just do it

OpenStudy (yash2651995):

\[(a+b)^{4}=a ^{4}+b ^{4}+4ab(a+b)^{2}-2(ab)^{2}\]

OpenStudy (shubhamsrg):

yep..same thing

OpenStudy (yash2651995):

ab=14 or 148..

OpenStudy (yash2651995):

a-b=5

OpenStudy (yash2651995):

a=7 b=2

OpenStudy (yash2651995):

or a=2 b=7